OpenStudy (cenaida):

Find the perimeter of the rectangle in simplest form. A rectangle with a width of 3 n squared + 7 n and a length of 5 n squared + 6 n + 3. A. 16n2 + 26n + 6 B. 15n2 + 42n + 3 C. 8n2 + 13n + 3

OpenStudy (anonymous):

Perimeter = 2(length) + 2(width) so substitue your length and width perimeter = 2(5n^2+6n) + 2(3n^2+7n) =10n^2+12n + 6n^2+14n combine like terms and add =16n^2 +26n+3

OpenStudy (anonymous):

+6 i mean

OpenStudy (cenaida):

thank you!!!

OpenStudy (anonymous):

np ^^)